A particle accelerate uniformly from a speed of 10m/s to 20m/s in 5 seconds.if it maintain this speed for 5 seconds before being brought to rest in another 5 seconds, calculate the total distance

1
Expert's answer
2021-11-01T19:24:58-0400

Using the kinematic formulas, find the distance travelled during the first part:


"d_1 = \dfrac{10m/s + 20m/s}{2}\cdot 5s = 75m"

During the second part:


"d_2 = 20m/s\cdot 5s = 100m"

And the third part:


"d_3 = \dfrac{0m/s + 20m/s}{5s} = 4m"

The total distance is:


"d = 75m + 100m + 4m = 179m"

Answer. 179 m.